Can you please help me with this problem? I'm confused. A square light switch plate is made which measures 4 inches on each side. A square hole measuring 1 inch on each side is left in the center. If the cost to construct the plate is 2.75 per square inch, what is the construction cost? (use A=s^2)
Answer by mananth(16949)

The area of the switch plate = area of square plate - the area of the square hole.
Area of sq. plate = 4*4 = 16 sq. inches
..
Area of square hole = 1*1=1 sq. in
Area of plate = 16-1= 15 sq. in
1 sq. inch cost $2.75
15 sq. in costs 2.75 * 15 =41.25